OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (4, 4) with 0° ≤ θ < 360°

OpenStudy (sidsiddhartha):

use x=rcos(theta) and y=rsin(theta) from here u can determine theta can u do it?

OpenStudy (anonymous):

not quite, what is r again?

OpenStudy (sidsiddhartha):

look in cartesian coordinate system there are mainly two coordinates which is (x,y) and in polar system the coordinates are expressed as (r,theta) caresian x is expressed as \[x=rcos(\theta)\] and cartesian y is expressed as \[y=rsin(\theta)\] make sense?

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

so \[\frac{ x }{ y }=rcos \theta/r \sin \theta\] so \[\tan \theta=y/x\]and then\[\theta=\tan^{-1} (\frac{ y }{ x })\] ok so far?

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

so now just put x=4 and y=4 so whats the value of theta?

OpenStudy (anonymous):

pi/4

OpenStudy (sidsiddhartha):

good so theta is evaluated now we need to calculate r

OpenStudy (sidsiddhartha):

for r -- find \[x^2+y^2\] can u do it?

OpenStudy (anonymous):

32

OpenStudy (sidsiddhartha):

yes so it would be equal to r^2 now ok with this?

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

\[x^2+y^2=r^2(\cos^2 \theta+\sin ^2 \theta)=r^2\] \[r=\sqrt{x^2+y^2}=\sqrt{32}=4\sqrt{2}\] got this?

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

so the polar coordinates are \[(4\sqrt{2},\frac{ \pi }{ 4 })\]

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

good :)

OpenStudy (anonymous):

But the answers are in degrees, so do we need to convert?

OpenStudy (sidsiddhartha):

yeah u can write (pi/4) equal to 45 degree

OpenStudy (anonymous):

ok, thanks

OpenStudy (sidsiddhartha):

yw!

OpenStudy (sidsiddhartha):

and u can click "best response" button if u feel that my help is enough ^_^